# Segment Anything Model for Medical Images?

### 12 Dec 2023 v5

@@ 1.2 SAM in Medical Image Analysis
Ji et al. (2023a) assessed the SAM under the Everything mode in segmenting the lesion regions in various anatomical structures (e.g., brain, lung, and liver) and modalities (computerized tomography (CT) and magnetic resonance imaging (MRI)). The experimental results indicated that SAM was relatively proficient in segmenting organ regions with clear boundaries but may struggle to accurately identify amorphous lesion regions. Another study then evaluated the performance of SAM in some healthcare subfields (optical disc and cup, polyp, and skin lesion segmentation) using both automatic Everything and two manual Prompt (points and boxes) strategies (Ji et al., 2023b). The authors found that SAM required substantial human prior knowledge (i.e., points) to obtain relatively accurate results on these tasks. Otherwise, SAM resulted in wrong segmentation, especially when no prompts are given. In the brain extraction task using MRI, Mohapatra et al. (2023) compared SAM with the brain extraction tool (BET) of the FMRIB Software Library. Quantitative results showed that the segmentation results of SAM were better than those of BET, demonstrating the potential of SAM for application in brain extraction tasks. Deng et al. (2023) assessed the performance of SAM in digital pathology segmentation tasks, including tumor, non-tumor tissue and cell nuclei segmentation on whole-slide imaging. The results suggested that SAM delivers outstanding segmentation results for large connected objects. Nevertheless, it may not consistently achieve satisfactory performance for dense instance object segmentation, even with prompts of all the target boxes or 20 points per image. Zhou et al. (2023) applied SAM to the polyp segmentation task using five benchmark datasets under the Everything setting. The results showed that although SAM can accurately segment the polyps in some cases, a large gap exists between SAM and the state-of-the-art methods. Additionally, Liu et al. (2023) equipped the 3D Slicer software (Pieper et al., 2004) with SAM to assist in the development, assessment, and utilization of SAM on medical images.
@@
Most recently, several studies tested the SAM on $\geq$10 public MIS datasets or tasks (He et al., 2023; Mazurowski et al., 2023; Ma and Wang, 2023; Wu et al., 2023). In (He et al., 2023), it was concluded that SAM’s zero-shot segmentation performance is considerably inferior to that of traditional deep learning-based methods. In Mazurowski et al. (2023), the authors evaluated SAM’s performance using different numbers of point prompts. They observed that as the number of points increases, the performance of SAM converges. They also noticed that SAM’s performance is 1) overall moderate and 2) extremely unstable across different datasets and cases. Ma and Wang (2023) validated that the original SAM may fail on lots of medical datasets with the mean DICE score of 58.52%. They then finetuned the SAM using medical images and found that the proposed MedSAM achieved a 22.51% improvement on DICE compared with the SAM. Wu et al. adopted the Adapter technique to finetune the SAM and enhance its medical ability. Experiments validated that their proposed Medical SAM Adapter can outperform the state-of-the-art (SOTA) MIS methods (e.g., nnUnet (Isensee et al., 2021)).
@@ 2 Dataset
Medical images have various modalities such as CT, MRI, ultrasound (US), and X-ray. There are large domain gaps between different modalities (Wang et al., 2020), and various modalities have their advantages in visualizing specific objects, including anatomical structures and lesions (Liu et al., 2023). To fully evaluate the generalization performance of SAM in MIS, we collected 53 public datasets and standardized them to construct the large COSMOS 1050K dataset. For the categorization system (e.g., modal categorization) of COSMOS 1050K, we referred to the official introduction of each public dataset and the recently-published study (Butoi et al., 2023) (see Table 1 for more details). Fig. 1 and Fig. 2 illustrate various imaging modalities and most of the clinical segmentation objects covered in the dataset, respectively. We describe the details of COSMOS 1050K in the following two aspects, including image collection and preprocessing specification.
@@
[b] Dataset Name Description Image Modalities AbdomenCT-1K (Ma et al., 2022) Liver, kidney, spleen and pancreas CT ACDC (Bernard et al., 2018) Left and right ventricle and left ventricular myocardium Cine-MRI AMOS 2022 (Ji et al., 2022) Abdominal multi-organ segmentation CT, MRI AutoLaparo (Wang et al., 2022) Integrated dataset with multiple image-based perception tasks Colonoscopy BrainPTM 2021 (Avital et al., 2019) white matter tracts T1W MRI, DW MRI BraTS20 (Menze et al., 2014; Bakas et al., 2017, 2018) Brain tumor T1W MRI, T2W MRI, T1-GD MRI, T2-FLAIR MRI CAMUS (Leclerc et al., 2019) Four-chamber and Apical two-chamber heart US CellSeg Challenge-NeurIPS 2022 (Ma et al., 2023) Cell segmentation Microscopy CHAOS (Kavur et al., 2021) Livers, kidneys and spleens CT, T1W MRI, T2W MRI CHASE-DB1 (Zhang et al., 2016) Retinal vessel segmentation Fundus Chest CT Segmentation (Dataset, a) Lungs, heart and trachea CT CRAG (Graham et al., 2019) Colorectal adenocarcinoma Histopathology crossMoDA (Shapey et al., 2019) Vestibular schwannoma T1W MRI CVC-ClinicDB (Bernal et al., 2015) Polyp Colonoscopy DRIVE (Liu et al., 2022b) Retinal vessel segmentation Fundus EndoTect 2020 (Hicks et al., 2021) Polyp Colonoscopy EPFL-EM (Lucchi et al., 2013) Mitochondria and synapses segmentation Electron Microscopy ETIS-Larib Polyp DB (Yoon et al., 2022) Polyp Colonoscopy FeTA (Payette et al., 2021) Seven tissues of the infant brain T2W MRI HaN-Seg (Podobnik et al., 2023) Healthy organs-at-risk near the head and neck CT I2CVB (Lemaître et al., 2015) Prostate T2W MRI iChallenge-AMD (Li et al., 2020) Optic disc and fovea Fundus iChallenge-PALM (Huazhu et al., 2019) Optic disc and lesions from pathological myopia patients Fundus IDRiD 2018 (Porwal et al., 2018) Optic disc, fovea and lesion segmentation Fundus iSeg 2019 (Sun et al., 2021) White matter, gray matter, and cerebrospinal fluid of infant brain T1W MRI, T2W MRI ISIC 2018 (Tschandl et al., 2018; Codella et al., 2018, 2019) Melanoma of skin Dermoscopy IXI (Dataset, c) Callosum T1W MRI KiPA22 (He et al., 2021, 2020; Shao et al., 2011, 2012) Kidney, tumor, renal vein and renal artery CT KiTS19 (Heller et al., 2020) Kidneys and tumors CT KiTS21 (Zhao et al., 2022) Kidneys, cysts, tumors, ureters, arteries and veins CT Kvasir-Instrumen (Jha et al., 2021) Gastrointestinal procedure instruments such as snares, balloons, etc. Colonoscopy Kvasir-SEG (Jha et al., 2020) Gastrointestinal polyp Colonoscopy LiVScar (Karim et al., 2016) Infarct segmentation in the left ventricle CMR LUNA16 (Setio et al., 2017) Lungs, heart and trachea CT M&Ms (Campello et al., 2021) Left and right ventricle and left ventricular myocardium CMR MALBCV-Abdomen (Zhou et al., 2019b) Abdominal multi-organ segmentation CT Montgomery County CXR Set (Jaeger et al., 2014) Lung X-ray MRSpineSeg (Pang et al., 2020) multi-class segmentation of vertebrae and intervertebral discs MRI MSD (Simpson et al., 2019) Large-scale collection of 10 Medical Segmentation Datasets CT, MRI, ADC MRI, T1W MRI, T2W MRI, T1-GD MRI, T2-FLAIR MRI NCI-ISBI 2013 (Li et al., 2013) Prostate (peripheral zone, central gland) T2W MRI PROMISE12 (Litjens et al., 2014) Prostate T2W MRI QUBIQ 2021 (Ji et al., 2021) Kidney, prostate, brain growth, and brain tumor CT, MRI, T1W MRI, T2W MRI, T1-GD MRI, T1-FLAIR MRI SIIM-ACR (Viniavskyi et al., 2020) Pneumothorax segmentation X-ray SKI10 (Lee et al., 2010) Cartilage and bone segmentation from knee data MRI SLIVER07 (Heimann et al., 2009) Liver CT ssTEM (Cardona et al., 2010) Neuronal structures Electron Microscopy STARE (Hoover et al., 2000) Retinal vessel segmentation Fundus TN-SCUI 2020 (Kang et al., 2022) Thyroid nodule US TotalSegmentator (Wasserthal et al., 2023) Multiple anatomic structures segmentation(27 organs, 59 bones, 10 muscles, and 8 vessels) CT VerSe19 (Sekuboyina et al., 2021) Spine or vertebral segmentation CT VerSe20 (Löffler et al., 2020; Liebl et al., 2021) Spine or vertebral segmentation CT Warwick-QU (Sirinukunwattana et al., 2017) Gland segmentation Histopathology WORD (Luo et al., 2022) Abdominal multi-organ segmentation CT 4C2021 C04 TLS01 (Dataset, b) Throat and hypopharynx cancer lesion area CT
@@ 2.1 Dataset Collection
Medical images cover a wide range of object types, such as brain organs and tumors (Simpson et al., 2019; Shapey et al., 2019; Bakas et al., 2018; Sun et al., 2021; Podobnik et al., 2023), lungs and hearts (Bernard et al., 2018; Jaeger et al., 2014; Viniavskyi et al., 2020; Setio et al., 2017), abdomen (Simpson et al., 2019; Zhou et al., 2019b; Zhao et al., 2022; Ma et al., 2022; Ji et al., 2022; Luo et al., 2022), spine (Sekuboyina et al., 2021; Löffler et al., 2020; Pang et al., 2020), cells (Lee et al., 2022; Ma et al., 2023), and polyps (Jha et al., 2020; Hicks et al., 2021). Table 1 presents a detailed list of the collected MIS datasets and Fig. 3 (a) shows the amount of each dataset after preprocessing. To be compatible with different modes of evaluating SAM, we employed the following exclusion criteria: 1) Exclude objects that are extremely small, such as the cochlea, shown in Fig. 4 (a), and ureter. This is due to the difficulty of automatically generating points or box prompts on extremely small objects. 2) Exclude objects in the 3D volume where their overall target became significantly separated as the slice was sequentially extracted, such as the intestine (seen in Fig. 4 (b)), mandible, and thyroid gland. We aim to avoid confusing the main object and generate unique boxes for each object. 3) Exclude objects with a relatively discrete overall structure, such as histopathological images of breast cancer (see Fig. 4 (c)), slices of lung trachea trees (see Fig. 4 (d)), renal arteries, and veins. Most of these objects are dispersed into multiple items in a 2D slice and embedded in other objects, resulting in a failure to sensibly employ SAM’s prompt mode on those objects to verify. According to the above criteria, COSMOS 1050K now comprises a total of 84 objects, with their numbers depicted in Fig. 3 (b). These objects are categorized only once in one image, without distinguishing locations or detailed divisions (e.g., ’Left Lung’ and ’Right Lung’ are grouped as ’Lung’, and various instruments are considered as ’Tool’). Further details can be found in the legend of Fig. 3. The histogram distributions of modalities and image resolutions are shown in Fig. 3 (c) and Fig. 3 (d), respectively. Given the significant variation of the same object across different modalities, including differences in gray distribution and texture features, we further divide them into 125 object-modality paired targets.
@@ 2.2 Dataset Preprocessing Specification
For 3D volumes, the whole procedure can be summarized as follows: 1) Extract slices along the main viewing plane owing to its higher resolution. In CT, it is usually a transverse plane, while in MRI, it may be a transverse plane, e.g., prostate, brain tumor, or a sagittal plane, e.g., spine and heart. 2) Retain slices with the sum of the pixel values of their labels greater than 50 for any 3D image and label volumes. This ensures that each slice has the corresponding correct label. 3) Normalize the extracted image intensities by min-max normalization: $I_{n}=255*(I-I_{min})/(I_{max}-I_{min})$, limiting the range to (0, 255). $I$ means the original extracted image, and $I_{n}$ represents the normalized image. $I_{min}$ and $I_{max}$ are the minimum and maximum intensity values of $I$. Simultaneously, we reset the pixel values of the mask according to the object’s category or location (e.g., left and right kidney have different pixel values). This is due to the fact that medical images may vary widely in the voxel or pixel values. Examples include MRI with an intensity range of (0, 800) and CT with an intensity range of (-2000, 2000), while other modalities may already be in the range of (0, 255) (Butoi et al., 2023). 4) Save images and labels in PNG format. For 4D data (N, W, H, D), we convert the data into N sets of 3D volume and then follow the 3D volume processing flow. Here, N represents the number of paired volumes within the 4D data.

### 17 Jan 2024 v7

@@ 1.2 SAM in Medical Image Analysis
Ji et al. (2023a) assessed the SAM under the Everything mode in segmenting the lesion regions in various anatomical structures (e.g., brain, lung, and liver) and modalities (computerized tomography (CT) and magnetic resonance imaging (MRI)). The experimental results indicated that SAM was relatively proficient in segmenting organ regions with clear boundaries but may struggle to accurately identify amorphous lesion regions. Another study then evaluated the performance of SAM in some healthcare subfields (optical disc and cup, polyp, and skin lesion segmentation) using both automatic Everything and two manual Prompt (points and boxes) strategies (Ji et al., 2023b). The authors found that SAM required substantial human prior knowledge (i.e., points) to obtain relatively accurate results on these tasks. Otherwise, SAM resulted in wrong segmentation, especially when no prompts are given. In the brain extraction task using MRI, Mohapatra et al. (2023) compared SAM with the brain extraction tool (BET) of the FMRIB Software Library. Quantitative results showed that the segmentation results of SAM were better than those of BET, demonstrating the potential of SAM for application in brain extraction tasks. Deng et al. (2023) assessed the performance of SAM in digital pathology segmentation tasks, including tumor, non-tumor tissue and cell nuclei segmentation on whole-slide imaging. The results suggested that SAM delivers outstanding segmentation results for large connected objects. Nevertheless, it may not consistently achieve satisfactory performance for dense instance object segmentation, even with prompts of all the target boxes or 20 points per image. Zhou et al. (2023) applied SAM to the polyp segmentation task using five benchmark datasets under the Everything setting. The results showed that although SAM can accurately segment the polyps in some cases, a large gap exists between SAM and the state-of-the-art methods. Additionally, Liu et al. (2023b) equipped the 3D Slicer software (Pieper et al., 2004) with SAM to assist in the development, assessment, and utilization of SAM on medical images.
@@
Most recently, several studies tested the SAM on $\geq$10 public MIS datasets or tasks (He et al., 2023; Mazurowski et al., 2023; Ma and Wang, 2023; Wu et al., 2023). In (He et al., 2023), it was concluded that SAM’s zero-shot segmentation performance is considerably inferior to that of traditional deep learning-based methods. In Mazurowski et al. (2023), the authors evaluated SAM’s performance using different numbers of point prompts. They observed that as the number of points increases, the performance of SAM converges. They also noticed that SAM’s performance is 1) overall moderate and 2) extremely unstable across different datasets and cases. Ma and Wang (2023) validated that the original SAM may fail on lots of medical datasets with the mean DICE score of 58.52%. They then finetuned the SAM using medical images and found that the proposed MedSAM achieved a 22.51% improvement on DICE compared with the SAM. Wu et al. (2023) adopted the Adapter technique to finetune the SAM and enhance its medical ability. Experiments validated that their proposed Medical SAM Adapter can outperform the state-of-the-art (SOTA) MIS methods (e.g., nnUnet (Isensee et al., 2021)).
@@ 2 Dataset
Medical images have various modalities such as CT, MRI, ultrasound (US), and X-ray. There are large domain gaps between different modalities (Wang et al., 2020), and various modalities have their advantages in visualizing specific objects, including anatomical structures and lesions (Liu et al., 2023b). To fully evaluate the generalization performance of SAM in MIS, we collected 53 public datasets and standardized them to construct the large COSMOS 1050K dataset. For the categorization system (e.g., modal categorization) of COSMOS 1050K, we referred to the official introduction of each public dataset and the recently-published study (Butoi* et al., 2023) (see Table 1 for more details). Fig. 1 and Fig. 2 illustrate various imaging modalities and most of the clinical segmentation objects covered in the dataset, respectively. We describe the details of COSMOS 1050K in the following two aspects, including image collection and preprocessing specification.
@@
[b] Dataset Name Description Image Modalities AbdomenCT-1K (Ma et al., 2021) Liver, kidney, spleen and pancreas CT ACDC (Bernard et al., 2018) Left and right ventricle and left ventricular myocardium Cine-MRI AMOS 2022 (Ji et al., 2022) Abdominal multi-organ segmentation CT, MRI AutoLaparo (Wang et al., 2022) Integrated dataset with multiple image-based perception tasks Colonoscopy BrainPTM 2021 (Avital et al., 2019; Nelkenbaum et al., 2020) white matter tracts T1W MRI, DW MRI BraTS20 (Menze et al., 2014; Bakas et al., 2017, 2018) Brain tumor T1W MRI, T2W MRI, T1-GD MRI, T2-FLAIR MRI CAMUS (Leclerc et al., 2019) Four-chamber and Apical two-chamber heart US CellSeg Challenge-NeurIPS 2022 (Ma et al., 2023) Cell segmentation Microscopy CHAOS (Kavur et al., 2021) Livers, kidneys and spleens CT, T1W MRI, T2W MRI CHASE-DB1 (Zhang et al., 2016) Retinal vessel segmentation Fundus Chest CT Segmentation (Dataset, a) Lungs, heart and trachea CT CRAG (Graham et al., 2019) Colorectal adenocarcinoma Histopathology crossMoDA (Shapey et al., 2019, 2021) Vestibular schwannoma T1W MRI CVC-ClinicDB (Bernal et al., 2015) Polyp Colonoscopy DRIVE (Liu et al., 2022) Retinal vessel segmentation Fundus EndoTect 2020 (Hicks et al., 2021) Polyp Colonoscopy EPFL-EM (Lucchi et al., 2013) Mitochondria and synapses segmentation Electron Microscopy ETIS-Larib Polyp DB (Bernal et al., 2017; Yoon et al., 2022) Polyp Colonoscopy FeTA (Payette et al., 2021) Seven tissues of the infant brain T2W MRI HaN-Seg (Podobnik et al., 2023) Healthy organs-at-risk near the head and neck CT I2CVB (Lemaître et al., 2015) Prostate T2W MRI iChallenge-AMD (Li et al., 2020) Optic disc and fovea Fundus iChallenge-PALM (Huazhu et al., 2019) Optic disc and lesions from pathological myopia patients Fundus IDRiD 2018 (Porwal et al., 2018) Optic disc, fovea and lesion segmentation Fundus iSeg 2019 (Sun et al., 2021) White matter, gray matter, and cerebrospinal fluid of infant brain T1W MRI, T2W MRI ISIC 2018 (Tschandl et al., 2018; Codella et al., 2018, 2019) Melanoma of skin Dermoscopy IXI (Dataset, c) Callosum T1W MRI KiPA22 (He et al., 2021, 2020; Shao et al., 2011, 2012) Kidney, tumor, renal vein and renal artery CT KiTS19 (Heller et al., 2020) Kidneys and tumors CT KiTS21 (Zhao et al., 2022) Kidneys, cysts, tumors, ureters, arteries and veins CT Kvasir-Instrumen (Jha et al., 2021) Gastrointestinal procedure instruments such as snares, balloons, etc. Colonoscopy Kvasir-SEG (Jha et al., 2020) Gastrointestinal polyp Colonoscopy LiVScar (Karim et al., 2016) Infarct segmentation in the left ventricle CMR LUNA16 (Setio et al., 2017) Lungs, heart and trachea CT M&Ms (Campello et al., 2021) Left and right ventricle and left ventricular myocardium CMR MALBCV-Abdomen (Landman et al., 2015) Abdominal multi-organ segmentation CT Montgomery County CXR Set (Jaeger et al., 2014) Lung X-ray MRSpineSeg (Pang et al., 2020) multi-class segmentation of vertebrae and intervertebral discs MRI MSD (Antonelli et al., 2022; Simpson et al., 2019) Large-scale collection of 10 Medical Segmentation Datasets CT, MRI, ADC MRI, T1W MRI, T2W MRI, T1-GD MRI, T2-FLAIR MRI NCI-ISBI 2013 (Li et al., 2013) Prostate (peripheral zone, central gland) T2W MRI PROMISE12 (Litjens et al., 2014) Prostate T2W MRI QUBIQ 2021 (Ji et al., 2021) Kidney, prostate, brain growth, and brain tumor CT, MRI, T1W MRI, T2W MRI, T1-GD MRI, T1-FLAIR MRI SIIM-ACR (Zawacki et al., 2019; Viniavskyi et al., 2020) Pneumothorax segmentation X-ray SKI10 (Lee et al., 2010) Cartilage and bone segmentation from knee data MRI SLIVER07 (Heimann et al., 2009) Liver CT ssTEM (Cardona et al., 2010) Neuronal structures Electron Microscopy STARE (Hoover et al., 2000; Hoover and Goldbaum, 2003) Retinal vessel segmentation Fundus TN-SCUI 2020 (Zhou et al., 2020) Thyroid nodule US TotalSegmentator (Wasserthal et al., 2023) Multiple anatomic structures segmentation(27 organs, 59 bones, 10 muscles, and 8 vessels) CT VerSe19 (Sekuboyina et al., 2021) Spine or vertebral segmentation CT VerSe20 (Löffler et al., 2020; Liebl et al., 2021) Spine or vertebral segmentation CT Warwick-QU (Sirinukunwattana et al., 2017) Gland segmentation Histopathology WORD (Luo et al., 2022) Abdominal multi-organ segmentation CT 4C2021 C04 TLS01 (Dataset, b) Throat and hypopharynx cancer lesion area CT
@@ 2.1 Dataset Collection
Medical images cover a wide range of object types, such as brain organs and tumors (Simpson et al., 2019; Shapey et al., 2019; Bakas et al., 2018; Sun et al., 2021; Podobnik et al., 2023), lungs and hearts (Bernard et al., 2018; Jaeger et al., 2014; Viniavskyi et al., 2020; Setio et al., 2017), abdomen (Simpson et al., 2019; Zhao et al., 2022; Ma et al., 2021; Ji et al., 2022; Luo et al., 2022), spine (Sekuboyina et al., 2021; Löffler et al., 2020; Pang et al., 2020), cells (Lee et al., 2023; Ma et al., 2023), and polyps (Jha et al., 2020; Hicks et al., 2021). Table 1 presents a detailed list of the collected MIS datasets and Fig. 3 (a) shows the amount of each dataset after preprocessing. To be compatible with different modes of evaluating SAM, we employed the following exclusion criteria: 1) Exclude objects that are extremely small, such as the cochlea, shown in Fig. 4 (a), and ureter. This is due to the difficulty of automatically generating points or box prompts on extremely small objects. 2) Exclude objects in the 3D volume where their overall target became significantly separated as the slice was sequentially extracted, such as the intestine (seen in Fig. 4 (b)), mandible, and thyroid gland. We aim to avoid confusing the main object and generate unique boxes for each object. 3) Exclude objects with a relatively discrete overall structure, such as histopathological images of breast cancer (see Fig. 4 (c)), slices of lung trachea trees (see Fig. 4 (d)), renal arteries, and veins. Most of these objects are dispersed into multiple items in a 2D slice and embedded in other objects, resulting in a failure to sensibly employ SAM’s prompt mode on those objects to verify. According to the above criteria, COSMOS 1050K now comprises a total of 84 objects, with their numbers depicted in Fig. 3 (b). These objects are categorized only once in one image, without distinguishing locations or detailed divisions (e.g., ’Left Lung’ and ’Right Lung’ are grouped as ’Lung’, and various instruments are considered as ’Tool’). Further details can be found in the legend of Fig. 3. The histogram distributions of modalities and image resolutions are shown in Fig. 3 (c) and Fig. 3 (d), respectively. Given the significant variation of the same object across different modalities, including differences in gray distribution and texture features, we further divide them into 125 object-modality paired targets.
@@ 2.2 Dataset Preprocessing Specification
For 3D volumes, the whole procedure can be summarized as follows: 1) Extract slices along the main viewing plane owing to its higher resolution. In CT, it is usually a transverse plane, while in MRI, it may be a transverse plane, e.g., prostate, brain tumor, or a sagittal plane, e.g., spine and heart. 2) Retain slices with the sum of the pixel values of their labels greater than 50 for any 3D image and label volumes. This ensures that each slice has the corresponding correct label. 3) Normalize the extracted image intensities by min-max normalization: $I_{n}=255*(I-I_{min})/(I_{max}-I_{min})$, limiting the range to (0, 255). $I$ means the original extracted image, and $I_{n}$ represents the normalized image. $I_{min}$ and $I_{max}$ are the minimum and maximum intensity values of $I$. Simultaneously, we reset the pixel values of the mask according to the object’s category or location (e.g., left and right kidney have different pixel values). This is due to the fact that medical images may vary widely in the voxel or pixel values. Examples include MRI with an intensity range of (0, 800) and CT with an intensity range of (-2000, 2000), while other modalities may already be in the range of (0, 255) (Butoi* et al., 2023). 4) Save images and labels in PNG format. For 4D data (N, W, H, D), we convert the data into N sets of 3D volume and then follow the 3D volume processing flow. Here, N represents the number of paired volumes within the 4D data.